OpenStudy (anonymous):

simplify the following. (x^-4*z^7)((2x^{2}y)/(z^-1))^-3

OpenStudy (anonymous):

Err \[(x^{-4}z^7)[{(2x^{2}y)\over(z^{-1})}]^{-3}\] This?

OpenStudy (anonymous):

\[ (x^{-4}*z^7)({\frac{2x^{2}y}{z^-1}})^{-3}\]

OpenStudy (anonymous):

Yeah, so what I wrote.

OpenStudy (anonymous):

yeah. z to the -1. z^(-1)

OpenStudy (anonymous):

So start by distributing the -3 exponent to the numerator & denominator.

OpenStudy (anonymous):

i got that. down. i have \[(x^{-4})\frac{z^3}{8x^6y^3}\]

OpenStudy (anonymous):

oops i mean. (x^-4*z^7)

OpenStudy (anonymous):

Good! Now combine the two factors

OpenStudy (anonymous):

Into one fractional expression I mean

OpenStudy (anonymous):

do i just multiply the (x^-4*z^7) by the numerator and that's all?

OpenStudy (anonymous):

And simplify, yes.

OpenStudy (anonymous):

ok so i got. \[\frac{z^{21}}{8xz^{12}x^6y^3}\]

OpenStudy (anonymous):

That's not right. I'm not sure how you got that actually? Can you explain?

OpenStudy (anonymous):

oops i messed up. should have added the exponents not multiplied. i see my mistake.

OpenStudy (anonymous):

When you multiply powers of the same base you do add their exponents.

OpenStudy (anonymous):

yeah i figured. thanks. appreciate it.